# Semiclassical treatment of bottomonium suppression and regeneration in $p+{\rm Pb}$ collisions

### v1

@@ II.1 Nuclear PDF (nPDF) Modifications
$\hskip-25.60747pt\times F_{i}^{p}(x_{1},\mu_{F}^{2},k_{T_{1}})~F_{j}^{p}(x_{2},\mu_{F}^{2},k_{T_{2}})~\hat{\sigma}_{ij}(\hat{s},\mu_{F}^{2},\mu_{R}^{2})\,$,
@@
where $i$, $j$ represent the incoming partons initiating $b\overline{b}$ production from processes $i+j=g+g$, $q+\overline{q}$, and, at next to leading order (NLO), $q(\overline{q})+g$, contributing to the partonic cross section $\hat{\sigma}_{ij}(\hat{s},\mu_{F}^{2},\mu_{R}^{2})$. The cross section depends on the quark mass, factorization scale $\mu_{F}$ and renormalization scale $\mu_{R}$. The NLO heavy-flavor cross section is obtained using the HVQMNR code [77]. The normalization factor $F_{C}$ is obtained by fitting the energy dependence of the summed $\Upsilon$(nnS) cross sections at $y=0$ measured through their decays to lepton pairs. (This summed branching ratio is taken into account because early $\Upsilon$ production measurements did not have sufficient mass resolution to differentiate among the three $\Upsilon$ peaks in the dilepton spectrum. The factor $F_{C}$ for individual $\Upsilon$ states can be obtained taking the branching ratios for specific states into account, along with the feed down to the inclusive $\Upsilon$(1S) from excited states.)
@@
The factors $F_{i,j}^{p}(x_{1,2},\mu_{F}^{2},k_{T_{1,2}})$ are the parton densities, assumed to factorize into the normal collinear parton densities and a $k_{T}$-dependent function as $F^{p}(x,\mu_{F}^{2},k_{T})=f^{p}(x,\mu_{F}^{2})G_{p}(k_{T})$. The CT10 PDFs [78], employed to determine $F_{C}$, are used for $f^{p}(x,\mu_{F}^{2})$. The function $G_{p}(k_{T})$, dependent on the intrinsic transverse momentum $k_{T}$, is assumed to be Gaussian [79], $G_{p}(k_{T})=\exp(-k_{T}^{2}/\langle k_{T}^{2}\rangle_{p})/{\pi\langle k_{T}^{2}\rangle_{p}}$.
@@
$\hskip-22.76219pt\times F_{i}^{p}(x_{1},\mu_{F}^{2},k_{T})~F_{j}^{A}(x_{2},\mu_{F}^{2},k_{T})~\hat{\sigma}_{ij}(\hat{s},\mu_{F}^{2},\mu_{R}^{2})\,.$
Here the parton densities in the nucleus, $F_{j}^{A}(x_{2},\mu_{F}^{2},k_{T})$, are taken to be the product of the nuclear modification factor $R_{j}(x_{2},\mu_{F}^{2},A)$ and the free parton density $F_{j}^{p}(x_{2},\mu_{F}^{2},k_{T})$ as,
$F_{j}^{A}(x_{2},\mu_{F}^{2},k_{T})$ $=$ $R_{j}(x_{2},\mu_{F}^{2},A)F_{j}^{p}(x_{2},\mu_{F}^{2},k_{T})\,.$ (3)
The NLO EPPS21 [25] nPDF set is used here to calculate the nPDF effect. The uncertainty on this result is obtained by calculating the cross sections using the central set and the 48 error sets (based on 24 fit parameters), added in quadrature, as also shown in Ref. [65]. Note that in other calculations in this approach, such as Ref. [24], momentum broadening in the nucleus is introduced in the Gaussian distribution, replacing $G_{p}(k_{T})$ with $G_{A}(k_{T})$. This modification is not included here to avoid doubling the effect of the broadening described in the following section.
@@ II.2 Energy loss and momentum broadening
where $y$ and $p_{T}$ are the bottomonium rapidity and transverse momentum respectively; the integral over $\varphi$ is defined as $\int_{\varphi}=\int_{0}^{2\pi}d\varphi/(2\pi)$; and $m_{T}=\sqrt{p_{T}^{2}+M_{\Upsilon}^{2}}$ is the transverse mass of the bottomonium state. The maximum rapidity shift is $\delta y_{\rm max}(y)=\text{min}(\ln 2,y_{\rm max}-y)$ where $y_{\rm max}=\ln(\sqrt{s_{NN}}/m_{T})$ is the maximum bottomonium rapidity in the $p+A$ center-of-momentum frame. The magnitude of the transverse momentum broadening is $\delta p_{T}=\sqrt{\ell^{2}_{A}-\ell^{2}_{p}}$ where $\ell_{A}$ and $\ell_{p}$ are the transverse momentum broadening due to the passage of the bottomonium state through a nucleus or a proton, respectively. The transverse momentum broadening acquired while the pair traverses the target nucleus $A$, as a function of the transport coefficient $\hat{q}$ and the effective path length $L_{A}$ through the target, is $\ell_{A}=\sqrt{\hat{q}L_{A}}$. The transport coefficient $\hat{q}$ is determined from fits to HERA data [80],
@@
The value of the parameter $\hat{q}_{0}=\hat{q}(x=10^{-2})$ employed here, $\hat{q}_{0}=0.075\;{\rm GeV}^{2}/{\rm fm}$, was determined in Ref. [28].
In Eq. (4), the quantity $\hat{\cal P}$ is the quenching weight that determines the probability for the energy loss and momentum broadening. It can be written in terms of the dilogarithm function ${\rm Li}_{2}$ as [29]
@@
where $N_{c}=3$ is the number of colors; the strong coupling constant, $\alpha_{s}$, is evaluated at the average momentum transfer; and $\delta x$ is the fractional loss, $\delta x=e^{\delta y}-1$ in Eq. (4). The numerator of the second term in Eq. (6), $\Lambda_{p}^{2}={\rm max}(\Lambda_{\mathrm{QCD}}^{2},\ell_{p}^{2})$, takes the value of either the square of momentum broadening in the proton, $\ell_{p}^{2}$, or the square of the QCD scale parameter, $\Lambda_{\rm QCD}^{2}$.
@@
The four-loop value of $\alpha_{s}$ is evaluated at the transverse momentum broadening scale $\ell_{A}$ with $\Lambda_{\rm QCD}=0.308$ GeV, resulting in $\alpha_{s}(1.5\;{\rm GeV})=0.326$, the value extracted from lattice QCD calculations of the static energy [81]. As in Ref. [65], we take average mass of $\Upsilon(1S)$, $\Upsilon(2S)$, and $\Upsilon(3S)$ states, $M=9.95$ GeV, for $M_{\Upsilon}$.
We adopt the following parametrization of the double differential prompt $\Upsilon$ cross section in Eq. (4[29]:
@@ II.3 Combined cold matter suppression
The factor $R^{\rm nPDF}_{pA}$ accounts for the suppression due to nPDF effects while $R^{\rm eloss,broad}_{pA}$ represents the suppression due to coherent energy loss and momentum broadening. We note here that the extraction of $\hat{q}_{0}$ in Ref. [28] from fixed-target $p+A$ data did not take nPDF effects into account. Likewise, global analyses of nPDFs do not include any energy loss effects in these extractions. As noted in Ref. [82], including $J/\psi$ data on nPDF extractions could reduce the nPDF uncertainties on the gluon distribution. However, the maximum gluon nPDF modification at low $x$ would depend on whether energy loss effects were considered or not.
@@ III QGP Quarkonium Suppression
The two transport coefficients governing the rates are $\Gamma_{\Upsilon}(T(\tau))$, the temperature and time-dependent inelastic reaction rate, and $N_{\Upsilon}^{\text{eq}}(\tau)$, the equilibrium limit on the number of $\Upsilon$ states. Here, $\Upsilon$ represents a generic bottomonium state. The rate equation can be decomposed into two parts: primordial suppression (first term in the right hand side of Eq. (10)) and regeneration (seconed term). In this section we will focus on primordial suppression, leaving the discussion of the regeneration to Sec. IV.
@@ III.1 Bottomonium reaction rates in a strongly-coupled QGP
Figure 3: Perturbative rates based on a in-medium $U$-potential for bottomonium binding, as functions of temperature at $p=0$. The solid lines represent the inelastic scattering rates while the dashed lines represent the gluo-dissociation rates for the 1S (blue), 1P (orange), 2S (green), and 3S (red) states.
@@
Figure 4: Perturbative bottomonium rates (inelastic scattering and gluon dissociation) assuming an in-medium $U$-potential, as functions of momentum for various temperatures. Each panel corresponds to a specific bottomonium state. The colors represent different temperatures, ranging from $T=170$ MeV (blue) to $T=700$ MeV (gray), as indicated in the legend.
@@
Figure 5: Nonperturbative bottomonium dissociation rates as a function of momentum over a range of temperatures. Each panel corresponds to a specific bottomonium state except the bottom right panel which shows twice the quark rates. The colors represent different temperatures, ranging from $T=170$ MeV (blue) to $T=700$ MeV (gray), as indicated in the legends. At high temperatures, the rates merge with the sum of the $b$ and $\overline{b}$ rates, shown in the bottom right panel.
The TAMU-NP rates are significantly larger than the TAMU-P rates, particularly for the excited states, as shown in Fig. 5. These rates generally decrease with increasing momentum, except for the $\Upsilon$(1S) rate at low temperature which is strongly suppressed at small $p$ due to the large 1S binding energy.
@@
Figure 6: Temperature profiles $T(\eta)$ at the transverse center of the medium (x=0,y=0)(x{=}0,\,y{=}0) for selected proper times $\tau$ in $p{+}\mathrm{Pb}$ collisions at $\sqrt{s_{NN}}=8.16$ and $5.02$ TeV from the aHydroQP simulation. The upper (a) and lower (b) panels show $\tau$-values relevant to the QTraj and TAMU calculations, respectively.
@@
Figure 6 compares the evolution of the longitudinal temperature profile from the aHydroQP simulation for $p{+}\mathrm{Pb}$ collisions at $\sqrt{s_{NN}}=5.02$ and $8.16$ TeV, shown at several relevant proper times. The hydrodynamic initialization time, $\tau=0.25$ fm/$c$, is shown in both panels. In the top panel, temperature profiles relevant for the OQS+pNRQCD calculation, labeled QTraj, are shown. The time $\tau=0.6$ fm/$c$ corresponds to when the $\Upsilon$ states begin to interact with the system in this approach while time at which the central temperature cools to $T_{\rm dec}=180$ MeV, $\tau=1.66$ fm/$c$, when the $\Upsilon$ interactions with the medium cease, are shown. In the bottom panel, the additional times are chosen to align with the formation times for the bottomonium states in the TAMU approach, indicated in Table 1, and the time at which these states cease interacting with the media at $T_{\rm dec}=170$ MeV.
@@
III.3 Effect of primordial suppression

### v2

@@ II.1 Nuclear PDF (nPDF) Modifications
$\hskip-22.76219pt\times F_{i}^{p}(x_{1},\mu_{F}^{2},k_{T_{1}})~F_{j}^{p}(x_{2},\mu_{F}^{2},k_{T_{2}})~\hat{\sigma}_{ij}(\hat{s},\mu_{F}^{2},\mu_{R}^{2})\,$,
where $i$, $j$ represent the incoming partons initiating $b\overline{b}$ production from processes $i+j=g+g$, $q+\overline{q}$, and, at next to leading order (NLO), $q(\overline{q})+g$, contributing to the partonic cross section $\hat{\sigma}_{ij}(\hat{s},\mu_{F}^{2},\mu_{R}^{2})$. The cross section depends on the quark mass, factorization scale $\mu_{F}$ and renormalization scale $\mu_{R}$. The NLO heavy-flavor cross section is obtained using the HVQMNR code [77]. The values for the bottom quark mass, $m$, and the scales $\mu_{F}$ and $\mu_{R}$, are determined from a fit to the total $b\overline{b}$ cross section at NLO: $(m,\mu_{F}/m_{T},\mu_{R}/m_{T})=(4.65\pm 0.09\,{\rm GeV},1.40^{+0.77}_{-0.59},1.10^{+0.22}_{-0.20})$. The scales are defined relative to the transverse mass of the $b\overline{b}$ pair, $\mu_{F,R}\propto m_{T}=\sqrt{m^{2}+p_{T}^{2}}$, where the $p_{T}$ is the transverse momentum of the $b\overline{b}$ pair, $p_{T}^{2}=0.5(p_{T,Q}^{2}+p_{T,{\overline{Q}}}^{2})$.
@@
The normalization factor $F_{C}$ is obtained by fitting the energy dependence of the summed $\Upsilon$(nnS) cross sections at $y=0$ measured through their decays to lepton pairs. A summed branching ratio was employed because early $\Upsilon$ measurements did not have sufficient mass resolution to differentiate among the three $\Upsilon$ peaks in the dilepton spectrum. The normalization factors of individual bottomonium states are determined based on the individual branching ratios to dileptons and feeddown from excited states.
@@
The parton densities in Eq. (II.1) include intrinsic $k_{T}$ broadening. These parton densities are assumed to factorize into the normal collinear parton densities and a $k_{T}$-dependent function, $F^{p}(x,\mu_{F}^{2},k_{T})=f^{p}(x,\mu_{F}^{2})G_{p}(k_{T})$. The CT10 proton parton distribution functions (PDFs) [78], used in the determination of $F_{C}$, are employed in the calculations of $f^{p}(x,\mu_{F}^{2})$.
The $k_{T}$ broadening is introduced to keep the $p_{T}$ distribution finite as $p_{T}\rightarrow 0$ because, at leading order (LO), the transverse momentum of the $Q\overline{Q}$ pair is zero. Broadening has typically been modeled by adding an intrinsic transverse momentum, $k_{T}$, to the parton densities. This broadening plays the role of low transverse momentum QCD resummation [79] without employing it directly. A $k_{T}$ broadening was initially added to the quarks in the initial state for Drell-Yan production [79] to create low $p_{T}$ lepton pairs because the Drell-Yan process also results in zero $p_{T}$ lepton pairs at leading order.
In the HVQMNR code, an intrinsic $k_{T}$ is added to each final state bottom quark, rather than to the initial state [77]. The $k_{T}$ broadening was introduced in Ref. [80] to compensate for the strong effect of fragmentation on fixed-target production of charm quarks. In the initial state, the intrinsic $k_{T}$ function multiplies the parton distribution functions for both hadrons, assuming the $x$ and $k_{T}$ dependencies factorize. At LO, there is no difference between adding a $k_{T}$ kick in the initial-state (the partons) or on the final-state (produced bottom quarks). However, at NLO, when there is a light parton in the final state, the correspondence is approximate. The difference between the two implementations is small if $\langle k_{T}^{2}\rangle\leq 2-3$ GeV2 [80].
A Gaussian distribution is employed for the broadening effect, $G_{p}(k_{T})=\exp(-k_{T}^{2}/\langle k_{T}^{2}\rangle_{p})/{\pi\langle k_{T}^{2}\rangle_{p}}$. It is applied by boosting the transverse momentum of the $b\overline{b}$ pair (plus light parton at NLO) to its rest frame from the longitudinal center-of-mass frame. The transverse momenta of the incident partons, $\vec{k}_{T\,1}$ and $\vec{k}_{T\,2}$, or, in this case, the final state $b$ and $\overline{b}$ quarks, are redistributed isotropically with unit modulus to satisfy momentum conservation. Once boosted back to the initial frame, the transverse momentum of the $b\overline{b}$ pair changes from $\vec{p}_{T}$ to $\vec{p}_{T}+\vec{k}_{T1}+\vec{k}_{T2}$ [81]. The rapidity distributions are independent of the intrinsic $k_{T}$.
The broadening effect decreases with increasing center of mass energy because the calculated average $p_{T}$ of the pair $b\overline{b}$ also increases with $\sqrt{s_{NN}}$. The value of $\langle k_{T}^{2}\rangle_{p}$ is assumed to increase with $\sqrt{s_{NN}}$, $\langle k_{T}^{2}\rangle_{p}=[1+(1/n)\ln(\sqrt{s_{NN}}/20)]$ GeV2, with $n=3$ for $\Upsilon$ production. Thus, at the LHC $p+A$ energies, $\langle k_{T}^{2}\rangle_{p}=2.84$ GeV2 and 3 GeV2 for $\sqrt{s_{NN}}=5.02$ and 8.16 TeV, respectively, within the range of applicability proposed in Ref. [80].
@@
$\hskip-42.67912pt\times F_{i}^{p}(x_{1},\mu_{F}^{2},k_{T\,1})~F_{j}^{A}(x_{2},\mu_{F}^{2},k_{T\,2})~\hat{\sigma}_{ij}(\hat{s},\mu_{F}^{2},\mu_{R}^{2})\,.$
Here the parton densities in the nucleus, $F_{j}^{A}(x_{2},\mu_{F}^{2},k_{T\,2})$, are taken to be the product of the nuclear modification factor $R_{j}(x_{2},\mu_{F}^{2},A)$ and the free parton density $F_{j}^{p}(x_{2},\mu_{F}^{2},k_{T})$ as,
$F_{j}^{A}(x_{2},\mu_{F}^{2},k_{T\,2})$ $=$ $R_{j}(x_{2},\mu_{F}^{2},A)F_{j}^{p}(x_{2},\mu_{F}^{2},k_{T\,2})\,.$ (3)
No additional $k_{T}$ broadening is included due to the presence of the nucleus. Instead, transverse momentum broadening is included, along with energy loss, in the following section.
The NLO EPPS21 [25] nPDF set is used here to calculate the nPDF effect. The uncertainty on this result is obtained by calculating the cross sections using the central set and the 48 error sets (based on 24 fit parameters), added in quadrature, as also shown in Ref. [65]. Note that in other calculations in this approach, such as Ref. [24], momentum broadening in the nucleus is introduced in the Gaussian distribution, replacing $G_{p}(k_{T})$ with $G_{A}(k_{T})$.
@@ II.2 Energy loss and momentum broadening
where $y$ and $p_{T}$ are the bottomonium rapidity and transverse momentum respectively; the integral over $\varphi$ is defined as $\int_{\varphi}=\int_{0}^{2\pi}d\varphi/(2\pi)$; and $m_{T}=\sqrt{p_{T}^{2}+M_{\Upsilon}^{2}}$ is the transverse mass of the bottomonium state where $M_{\Upsilon}=9.95$ GeV is the average mass of $\Upsilon$(1S), $\Upsilon$(2S), and $\Upsilon$(3S) states [65]. The maximum rapidity shift is $\delta y_{\rm max}(y)=\text{min}(\ln 2,y_{\rm max}-y)$ where $y_{\rm max}=\ln(\sqrt{s_{NN}}/m_{T})$ is the maximum bottomonium rapidity in the $p+A$ center-of-momentum frame. The magnitude of the transverse momentum broadening is $\delta p_{T}=\sqrt{\ell^{2}_{A}-\ell^{2}_{p}}$ where $\ell_{A}$ and $\ell_{p}$ are the transverse momentum broadening due to the passage of the bottomonium state through a nucleus or a proton, respectively. The transverse momentum broadening acquired while the pair traverses the target nucleus $A$, as a function of the transport coefficient $\hat{q}$ and the effective path length $L_{A}$ through the target, is $\ell_{A}=\sqrt{\hat{q}L_{A}}$. Similarly, $\ell_{p}=\sqrt{\hat{q}L_{p}}$ with $L_{\rm Pb}=10.41$ fm and $L_{p}=1.5$ fm [82]. The transport coefficient $\hat{q}$ is determined from fits to HERA data [83],
@@
The value of the parameter $\hat{q}_{0}=\hat{q}(x=10^{-2})$, $\hat{q}_{0}=0.075\;{\rm GeV}^{2}/{\rm fm}$, was determined in Ref. [28].
In Eq. (4), $\hat{\cal P}$ is the quenching weight that determines the probability for the energy loss and momentum broadening. It can be written in terms of the dilogarithm function ${\rm Li}_{2}$ as [29]
@@
where $N_{c}=3$ is the number of colors and $\delta x$ is the fractional loss, $\delta x=e^{\delta y}-1$ in Eq. (4). The numerator of the second term in Eq. (6), $\Lambda_{p}^{2}={\rm max}(\Lambda_{\mathrm{QCD}}^{2},\ell_{p}^{2})$, takes the value of either the square of momentum broadening in the proton, $\ell_{p}^{2}$, or the square of the QCD scale parameter, $\Lambda_{\rm QCD}^{2}$.
@@
As in Ref. [65], we employ the four-loop value of $\alpha_{s}$ in Eq. (6), evaluated at $\Lambda_{\rm QCD}=\delta p_{T}$ with $\Lambda_{\rm QCD}=0.308$ GeV to give $\alpha_{s}(1.5\;{\rm GeV})=0.326$, the value extracted from lattice QCD calculations of the static energy [84]. The original formulation of the quenching weights [28, 31] assumed the fixed coupling $\alpha_{s}=0.5$. We have determined that the difference between these choices is small and has no effect on the results of the calculation since $R_{pA}$ due to energy loss begins to differ in the two assumptions at $|y|>3$ and never more than 10% in the measured range, considerably smaller than the uncertainty band for the nPDF contribution.
We adopt the parametrization of the double differential prompt $\Upsilon$ cross section in Eq. (4) [29],
@@ II.3 Combined cold matter suppression
The factor $R^{\rm nPDF}_{pA}$ accounts for the suppression due to nPDF effects while $R^{\rm eloss,broad}_{pA}$ represents the suppression due to coherent energy loss and momentum broadening.
The results shown in Fig. 1 do not include any mass or scale variations in Eqs. (II.1) and (II.1) because we assume the mass and scale variations between the $p+p$ and $p+A$ calculations to be correlated [74]. Of course, taken separately, the mass and scale uncertainties are larger than the nPDF effects in quadrature.
We briefly comment on the combination of nPDFs with energy loss effects. The nPDFs are extracted from a variety of QCD processes, both in lepton-nucleus and in proton-nucleus collisions. The gluon distribution in the nucleus is typically only constrained by sum rules and QCD evolution. However, studies have shown that the production of open-charm mesons and charmonia in $p+A$ collisions at LHC energies provides a significant constraint on the small $x$ gluon distribution, see e.g. [25]. Because these processes are also sensitive to energy loss effects [28, 29, 85], these effects are thus absorbed into the nPDFs. However, not including the charm measurements would weaken the gluon nPDF constraints.
A recent paper [86] incorporated LO $J/\psi$ production in $p+A$ collisions in the CEM in calculations with several nPDF sets. They found the gluon nPDF ratio, $R_{g}$, using the $J/\psi$ data with energy loss included is higher than when energy loss is not included. See also Ref. [85]. Thus not including energy loss effects may lead to a lower central value of the gluon nPDF, most important for quarkonium production at the LHC. There is some kinematics mismatch in the LO approach since the nPDFs are extracted at next-to-leading order and the $J/\psi$ transverse momentum cannot be incorporated at LO.
We note also that the gluon ratio for EPPS21, which included LHC $D$ meson data in the analysis [25], is quite similar to that of EPPS16 [87] which did not. It is also worth mentioning that the value of $\hat{q}_{0}$ in Ref. [28], used here, did not incorporate nPDF. However, when saturation effects were included with the same $\hat{q}_{0}$, the value of $R_{pA}$ is decreased for $J/\psi$ production while the effect on $\Upsilon$ production was negligible.
@@ III Quarkonium Suppression in the QGP
The two transport coefficients governing the rates are $\Gamma_{\Upsilon}(T(\tau))$, the temperature and time-dependent inelastic reaction rate, and $N_{\Upsilon}^{\text{eq}}(\tau)$, the equilibrium limit on the number of $\Upsilon$ states. Here, $\Upsilon$ represents a generic bottomonium state. The rate equation can be decomposed into two parts: primordial suppression (first term in the right hand side of Eq. (10)) and regeneration (second term). In this section we will focus on primordial suppression, leaving the discussion of the regeneration to Sec. IV.
@@ III.1 Bottomonium reaction rates in a strongly-coupled QGP
Figure 3: Perturbative rates based on a in-medium $U$-potential for bottomonium binding, as functions of temperature at $p=0$ [15]. The solid lines represent the inelastic scattering rates while the dashed lines represent the gluo-dissociation rates for the 1S (blue), 1P (orange), 2S (green), and 3S (red) states.
@@
Figure 4: Perturbative bottomonium rates (inelastic scattering and gluon dissociation) assuming an in-medium $U$-potential, as functions of momentum for various temperatures [15]. Each panel corresponds to a specific bottomonium state. The colors represent different temperatures, ranging from $T=170$ MeV (blue) to $T=700$ MeV (gray), as indicated in the legend.
@@
Figure 5: Nonperturbative bottomonium dissociation rates obtained in the TAMU-NP framework, shown as functions of temperature at $p=0$ [67]. The rates are extracted from the pole structure of the in-medium $T$-matrix in the complex energy plane, using a heavy-quark potential constrained by nonperturbative lattice-QCD Wilson line correlators. The solid lines show the inelastic dissociation rates for the $\Upsilon$(1S) (blue), $\chi_{b}$(1P), $\Upsilon$(2S) (orange), $\chi_{b}$(2P) (purple), and $\Upsilon$(3S)(red) states.
Figure 6: Nonperturbative bottomonium dissociation rates as a function of momentum over a range of temperatures [67]. Each panel corresponds to a specific bottomonium state except the bottom right panel which shows twice the quark rates. The colors represent different temperatures, ranging from $T=170$ MeV (blue) to $T=700$ MeV (gray), as indicated in the legends. At high temperatures, the rates merge with the sum of the $b$ and $\overline{b}$ rates, shown in the bottom right panel.
The TAMU-NP rates are significantly larger than the TAMU-P rates, particularly for the excited states, as shown in Fig. 5 as a function of temperature for $p=0$ and in Fig. 6 as a function of momentum for a range of temperatures. These rates generally decrease with increasing momentum, except for the $\Upsilon$(1S) rate at low temperature which is strongly suppressed at small $p$ due to the large 1S binding energy.
@@
Figure 7: Temperature profiles $T(\eta)$ at the transverse center of the medium (x=0,y=0)(x{=}0,\,y{=}0) for selected proper times $\tau$ in $p{+}\mathrm{Pb}$ collisions at $\sqrt{s_{NN}}=8.16$ and $5.02$ TeV from the aHydroQP simulation. The upper (a) and lower (b) panels emphasize times $\tau$ (in fm) relevant to the QTraj and TAMU calculations, respectively.
@@
Figure 7 compares the evolution of the longitudinal temperature profile from the aHydroQP simulation for $p{+}\mathrm{Pb}$ collisions at $\sqrt{s_{NN}}=5.02$ and $8.16$ TeV, for several different proper times relevant in the OQS+pNRQCD and TAMU approaches. The hydrodynamic initialization time, $\tau=0.25$ fm/$c$, is shown in both panels. In the top panel, temperature profiles relevant for the OQS+pNRQCD calculation, labeled QTraj, are shown. The time $\tau=0.6$ fm/$c$ corresponds to when the $\Upsilon$ states begin to interact with the system in this approach. The time at which the central temperature cools to $T_{\rm dec}=180$ MeV, $\tau=1.66$ fm/$c$, when the $\Upsilon$ interactions with the medium cease, is also given. In the bottom panel, the times $\tau=0.34$, 0.68, and 0.97 fm/c/c are representative of the bottomonium formation times in the TAMU approach, given in Table 1. The value $\tau=0.34$ fm/c/c is an average between the 1P and 2S formation times, 0.66 is the 2P formation time while 0.97 is that of the 3S. The time at which these states cease interacting with the medium at $T_{\rm dec}=170$ MeV, $\tau=1.85$ fm is also shown.
@@
III.3 Effect of primordial suppression alone